OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

x^2 +2

OpenStudy (anonymous):

wait what can you do the full eqaution

OpenStudy (anonymous):

ok i got this so this is g(3)=3^2+2 so 3^2=9 so you got g(3)=9+2.....9+2=11 so g(3)=11

OpenStudy (anonymous):

thanks babe

OpenStudy (anonymous):

your welcome babe